[Hallc_running] [Revised Logentry] Run Status

jones at jlab.org jones at jlab.org
Mon Jan 20 08:50:01 EST 2020

Logentry Text:
We had a good Day shift 6.7 out of 8.
During swing there was a problem because of a failed vacuum gauge causing 8A/8R valves to close and 9S BLM trips when C beam was on. 
Swing was only 2.3 out of 8 hours
Reza Kazimi came and from midnight to 2 adjusted the injector, and got 30 uA back to C.
Owl was 5.6 out of 8 hours.

Jay has a summary of yesterday's machine performance at https://logbooks.jlab.org/entry/3767587

With Hall A off the charge asymmetry feedback was moved to Hall C.
See log entry https://logbooks.jlab.org/entry/3767296 about how to recovery if there are problems.

We will keep doing NMR measurements every 5-6 hours. 

We will do a flip of the insertable half wave plate this morning. We will leave in this state until Tuesday DAY shift.

Tuesday DAY shift will be a accelerator work in the NL on  1L20  heat load problems. Jay estimates that the work and recovery
will take most of DAY shift.

We will develop a plan on what to do during this downtime on Tuesday.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3767599
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200120/a1962f38/attachment.html>

More information about the Hallc_running mailing list